OpenStudy (anonymous):

*** Will give you a medal *** Describe the vertical asymptote(s) and hole(s) for the graph of y = (x-5)/(x^2+4x+3)

OpenStudy (anonymous):

@kropot72

OpenStudy (anonymous):

\[\lim_{x \rightarrow \infty}\frac{x-5}{x^2+4x+3}=0\] Therefore y=0 is an retricemptote. \[y=\frac{x-5}{(x+3)(x+1)}\] What value of x makes that equation become undefined? Undefined as in, what value of x makes the denominator become 0?

OpenStudy (anonymous):

asymptote*

OpenStudy (anonymous):

0?

OpenStudy (anonymous):

asymptotes: x = –3, –1 and no holes. asymptote: x = –3 and hole: x = –5 asymptotes: x = –3, –1 and hole: x = –5 asymptote: x = –5 and hole: x = –3 These are my options

OpenStudy (anonymous):

@jim_thompson5910

OpenStudy (anonymous):

Dude...solve this quadratic for x. \[(x+3)(x+1)=0\] What values of x do you get from that?